OpenStudy (clamin):

HELP!! MEDAL!! A high school is installing a brand new regulation soccer field measuring 100 yards by 60 yards. In order to meet zoning regulations, the field must have a 6-inch layer of gravel under the topsoil to provide adequate drainage.The gravel supplier charges either $12.50 per cubic yard of gravel or a flat rate of $95 for a full truck carrying 9 cubic yards of gravel. Question 1: Find the area of the soccer field. Question 2:Find the volume of gravel necessary to cover the soccer field. Question 3:How much money should the school budget for gravel to install its new field?

OpenStudy (butterflydreamer):

okayy. Can you do question 1 ? :)

OpenStudy (butterflydreamer):

Question 1 is basic area... of a rectangle. Area of a rectangle = length * width. We are given that the measurements are 100 yards by 60 yards. Therefore Area = 100 * 60 = ? Question 2 : We are trying to find the volume of the gravel. The first step is to CONVERT 6 inches INTO yards 6 inches = 0.167 yards. Now that's done, to find the volume we do : Volume = length * width * height. Volume = 100 * 60 * 0.167 = ? cubic yards Question 3: Once you have the answer to question 2, we use the info given to us "The gravel supplier charges either $12.50 per cubic yard of gravel or a flat rate of $95 for a full truck carrying 9 cubic yards of gravel."